# Mixture of Cluster-conditional LoRA Experts for Vision-language Instruction Tuning

### 18 Jan 2024 v2

@@ Abstract
Instruction tuning of the Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflicts for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve the generalization capabilities of MoCLE for novel instructions. Extensive experiments on 10 zero-shot tasks demonstrate the effectiveness of MoCLE.
@@ 4.1 Settings and Implementation Details
Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 44.2 64.3 30.3 59.4 41.5 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 42.7 65.3 29.5 62.0 41.4 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 45.3(+2.6) 65.6(+0.3) 30.4(+0.9) 63.9(+1.9) 42.6(+1.2) 53.2(+1.9)
@@ 4.2 Datasets and Instructions
For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], and iVQA [57]. InstructBLIP originally uses the NoCaps dataset [1] for the image captioning task, which is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used. In addition, we do not utilize the MSRVTT-QA dataset [56] due to the insufficient computational resources.
@@ 5 Conclusions
In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of the proposed MoCLE method.
@@ Appendix H Case Studies
In Table 11, the model is queried to recognize the character in the image. InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training. With MoCLE, the model can give correct results.
@@
GQA, VizWiz, iVQA, MSVD Question: {} Short answer:

### 08 Feb 2024 v3

@@ Abstract
Instruction tuning of the Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflicts for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve the generalization capabilities of MoCLE for novel instructions. Extensive experiments on 12 zero-shot tasks demonstrate the effectiveness of MoCLE.
@@ 4.1 Settings and Implementation Details
Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA MSRVTT QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 17.4 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 9.2 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 22.1 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 46.0 64.3 33.7 59.4 41.5 22.7 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 46.3 65.3 34.5 62.0 41.4 23.0 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 46.9(+0.6) 65.6(+0.3) 34.5(+0.0) 63.9(+1.9) 42.6(+1.2) 24.4(+1.4) 53.2(+1.9)
@@ 4.2 Datasets and Instructions
For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], MSRVTT-QA [56] and iVQA [57]. InstructBLIP originally uses the NoCaps dataset [1] for the image captioning task, which is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used.
@@ 5 Conclusions
In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on 12 unseen tasks show the effectiveness of the proposed MoCLE method.
@@ Appendix H Case Studies
In Table G, the model is queried to recognize the character in the image. InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training. With MoCLE, the model can give correct results.
@@
GQA, VizWiz, iVQA, MSVD, MSRVTT Question: {} Short answer: